Find a polynomial for the perimeter and for the area. c+8, c
Answer by stanbon(75887)

Find a polynomial for the perimeter and for the area. c+8, c.
---
If the dimensions are c and c+8,
Perimeter = 2(c+c+8) = 2(2c+8) = 4c+16 units
---
Area = c(c+8) = c^2+8c sq. units
====================
Cheers,
Stan H.